# Complexity of Hofer’s geometry in higher dimensional manifolds

### v1

@@ 3.4. Loop rotation operator and $\mathbb{Z}_{k}$ spectral spread
The following proposition, generalizing ideas from [PS, Propositions 4.2 and 4.3], provides the key estimate.
@@
Let $H:S^{1}\times M\to\mathbb{R}$ be a Hamiltonian on $M$. Suppose $H^{(k)}(t,x)=kH(kt,x)$ is non-degenerate and $\alpha\in\pi_{0}(\mathcal{L}M)$ is a primitive element (i.e. it is not a power of another element). For $r\in\mathbb{Z}$, suppose that all generators $p$ of $\mathrm{CF}_{*}(H^{(k)})_{\alpha}$ with index $r$ satisfy
$\min_{q\in\partial^{-1}(p)}\left(\mathcal{A}_{H^{(k)}}(q)-\mathcal{A}_{H^{(k)}}(p)\right)\geq D>0.$
Then the Hofer distance is bounded below by $d_{H}(\phi_{H},\mathrm{Ham}^{k}(M))\geq D$.
@@
The proof follows the same idea as [PS, Propositions 4.2 and 4.3].
The key object is the loop rotation operator
@@
which induces a map on the filtered Floer complex $\mathrm{CF}_{*}(H^{(k)})_{\alpha}$. By [PS, Lemma 3.1], this map agrees with the action of $(\phi_{H})_{*}$ on the filtered Floer homology.
Since $\alpha$ is primitive, for a generator $x$, its orbit $R_{k}^{i}(x):1\leq i\leq k$ under the $\mathbb{Z}_{k}$-action generated by $R_{k}$ are all distinct.
Consider the operator $S_{k}=R_{k}-\mathrm{id}$ on the Floer complex. By our assumption on the action difference, for any $\epsilon>0$, there exist an action window $(a,b)$ and $d>D-2\epsilon$ such that
$j_{d}\circ S_{k}:\mathrm{HF}_{*}(H^{(k)})_{\alpha}^{(a,b)}\to\mathrm{HF}_{*}(H^{(k)})_{\alpha}^{(a+d,b+d)}$
is non-zero. Then we use [PS, Proposition 4.2(ii)] and let $\epsilon\to 0$ to prove the result. ∎
@@ 4. Generalized linked twist map on plumbing of tori
Furthermore, if $w=(a^{k_{1}}b^{k_{2}})^{m}$, then the orbit class $\gamma(N,w)$ can be chosen to be primitive.
@@ 4.1. Proof of the main theorems
For $\geq 2, kk\in\mathbb{N}$, take $f_{N}=\tau(N,ab)$ for $ab\in\mathbb{F}_{2}=\langle a,b\rangle$. Then the $k$-periodic points of $f_{N}$ are exactly the fixed points of $\tau(N,(ab)^{k})$. By Proposition 4.1, there exists an index $r$ and a primitive orbit class $\gamma$ such that the action difference between any fixed point $p$ of $\tau(N,(ab)^{k})$ and any $q\in\partial^{-1}(p)$ is at least $\delta N$.
Therefore, by Proposition 3.2, we have $d_{H}(f_{N},\mathrm{Ham}^{k}(M))\geq\delta N$, which implies $d_{H}(f_{N},\mathrm{Ham}^{k}(M))\to\infty$ as $N\to\infty$. ∎
@@ 5. Conley–Zehnder index
From Proposition 5.1, there exists a unique fixed point $p=(v_{0},x_{0})$ of $\tau(N,w)$ with maximal Conley–Zehnder index $r$, such that its intermediate points $(v_{i},x_{i})$ given in Equation 4.4 satisfy $x_{i}\in B_{i}^{-\mathrm{sign}(k_{2i-1})}$ and $v_{i}\in C_{i}^{-\mathrm{sign}(k_{2i})}$. Let $q\in\partial^{-1}(p)$. Then the Conley–Zehnder index of $q$ is r1r-1, and the signs $\sigma_{i}$ and $\xi_{i}$ for $p$ and $q$ are the same except for one entry, which we assume without loss of generality to be $\sigma_{j}$.
@@
By definition, the fixed points in the orbit of $\tau(N,w_{1})$ have the same index and action. Then the same argument as above shows that there is a unique set of fixed points with maximal index, with action difference larger than $\frac{N\epsilon^{2}}{10\max_{j}|k_{j}|}+O(1)$.

### v2

@@ 3.4. Loop rotation operator and $\mathbb{Z}_{k}$ spectral spread
The following proposition, which is a corollary of Theorem 4.22 in [PS], provides the key estimate.
@@
Let $H:S^{1}\times M\to\mathbb{R}$ be a Hamiltonian on $M$. Suppose $H^{(k)}(t,x)=kH(kt,x)$ is non-degenerate and $\alpha\in\pi_{0}(\mathcal{L}M)$ is a primitive element (i.e. it is not a power of another element). Let $k$ be a prime number. For $r\in\mathbb{Z}$, suppose that there exists exactly $k$ many generators $p_{i},1\leq i\leq k$ of $\mathrm{CF}_{*}(H^{(k)})_{\alpha}$ with index $r$, and they satisfy
$\min_{q\in\partial^{-1}(p_{i})}\left(\mathcal{A}_{H^{(k)}}(q)-\mathcal{A}_{H^{(k)}}(p_{i})\right)\geq D>0.$
Then the Hofer distance is bounded below by $d_{H}(\phi_{H},\mathrm{Ham}^{k}(M))\geq\frac{1}{4k}D$.
@@
We consider the loop rotation operator
@@
which induces $\mathbb{Z}_{k}$-action on the filtered Floer complex $\mathrm{CF}_{*}(H^{(k)})_{\alpha}$. By [PS, Lemma 3.1], this map agrees with the action of $(\phi_{H})_{*}$ on the filtered Floer homology.
Since $\alpha$ is primitive, for a generator $x$, its orbit $R_{k}^{i}(x):1\leq i\leq k$ under the $\mathbb{Z}_{k}$-action generated by $R_{k}$ are all distinct, therefore, by our assumption the generators of $\mathrm{CF}_{*}(H^{(k)})_{\alpha}$ with index $r$ are exactly the set $\{R_{k}^{i}(x)\}$.
This shows that for any unitary root $\xi\in\mathbb{Z}_{k}$, there exists exactly one element of index $r$ in the eigenspace of $\xi$ under the $\mathbb{Z}_{k}$ action of $\mathrm{CF}_{*}(H^{(k)})_{\alpha}$, which does not vanish on an interval of length $D$.
Therefore, by Theorem 4.22 of [PS], we conclude that $d_{H}(\phi_{H},\mathrm{Ham}^{k}(M))\geq\frac{1}{4k}(D-2\epsilon)$ for any $\epsilon>0$, let $\epsilon\to 0$ yeilds the desired result. ∎
@@ 4. Generalized linked twist map on plumbing of tori
Furthermore, if $w=(a^{k_{1}}b^{k_{2}})^{m}$, then $\gamma(N,w)$ can be chosen to be primitive, and there exist exactly $m$ fixed points of $\tau(N,w)$ of index $r$ in the class $\gamma(N,w)$.
@@ 4.1. Proof of the main theorems
For a prime number $\geq 2, kk\in\mathbb{N}$, take $f_{N}=\tau(N,ab)$ for $ab\in\mathbb{F}_{2}=\langle a,b\rangle$. Then the $k$-periodic points of $f_{N}$ are exactly the fixed points of $\tau(N,(ab)^{k})$. By Proposition 4.1, there exists an index $r$ and a primitive orbit class $\gamma$ such that there exists exactly $k$ fixed points of $\tau(N,(ab)^{k})$ of index $r$ and the action difference between any fixed point $p$ of index $r$ and any $q\in\partial^{-1}(p)$ is at least $\delta N$.
Therefore, by Proposition 3.2, we have $d_{H}(f_{N},\mathrm{Ham}^{k}(M))\geq\frac{1}{4k}\delta N$, which implies that $d_{H}(f_{N},\mathrm{Ham}^{k}(M))\to\infty$ as $N\to\infty$.
If $k$ is not a prime integer, then $\mathrm{Ham}^{k}(M)\subset\mathrm{Ham}^{k^{\prime}}(M)$ for some prime number $k^{\prime}$, therefore, we still have $d_{H}(f_{N},\mathrm{Ham}^{k}(M))\to\infty$ as $N\to\infty$. ∎
@@ 5. Conley–Zehnder index
From Proposition 5.1, there exists a unique fixed point $p=(v_{0},x_{0})$ of $\tau(N,w)$ with minimum Conley–Zehnder index $r$, such that its intermediate points $(v_{i},x_{i})$ given in Equation 4.4 satisfy $x_{i}\in B_{i}^{-\mathrm{sign}(k_{2i-1})}$ and $v_{i}\in C_{i}^{-\mathrm{sign}(k_{2i})}$. Let $q\in\partial^{-1}(p)$. Then the Conley–Zehnder index of $q$ is r1r-1, and the signs $\sigma_{i}$ and $\xi_{i}$ for $p$ and $q$ are the same except for one entry, which we assume without loss of generality to be $\sigma_{j}$.
@@
By definition, the fixed points in the orbit of $\tau(N,w_{1})$ have the same index and action. Then by Proposition 5.1 the same argument as above shows that there is a unique set of fixed points with minimum index, with action difference larger than $\frac{N\epsilon^{2}}{10\max_{j}|k_{j}|}+O(1)$.